Find the vertex and the x-intercepts of the parabola: Show Steps
y = 3x^2+5x–7

Answer by addingup(3677)

Vertex: (-5/6, -109/12)~~(-0.833333, -9.08333)
Intercepts:
y = -7
x = 1/6 (-5-sqrt(109)), x = 1/6 (sqrt(109)-5)